Plz specify the degree & order of the following eq with reasons (keeping in mind that sinx,cosx etc. can be expanded polynomially -> a)d^2y/dx^2+sin(dy/dx)=0 b)d2y/dx2+sin(d3y/dx3)=0 c)(d2y/dx2)^3 + sin(dy/dx)= 0 d)d2y/dx2+sin{(d3y/dx3)^4}=0
what will be the ans if there is other transcendental Fn in place of sin?
Plz say me urgently.

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

 

Dear Borah, 

a)d2y/dx2+sin(dy/dx)=0, based on definitions of order and degree, order=2,degree=1

b)d2y/dx2 +sin(d3y/dx3)=0, d3y/dx3+arcsin(d2y/dx2)=0,order=3 and degree=1

c)order=2 and degree=3

d)(d3y/dx3)4 +arcsin(d2y/dx2)=0, order=3 and degree=4

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras